# Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs

### v1

@@ 5.2. End-to-end Evaluation
5.3. Ablation Studies
Next, we assess the effectiveness of the proposed techniques in Memo. All experiments in ablation studies are conducted by training the 7B model on 8 GPUs, keeping the parallelism configuration fixed at a TP size of 4 and a CP size of 2.
5.3.1. Effectiveness of Memory Planning
Additionally, the frequent GPU memory reorganization process further impairs training efficiency. By employing memory planning, the fragmentation issue can be minimized, providing more memory for longer context training. Getting rid of GPU memory reorganization, memory planning brings an average of $1.51\times$ MFU when facing the same context length.
5.3.2. Effectiveness of Token-wise Recomputation
For token-wise recomputation and swapping, we compare Memo and its variants, one with full recomputation and another with full swapping. The results are shown in the last two rows of Table 4. When training with appropriate sequence length, which is 256k in this scenario, the computation time of one transformer layer can fully overlap with the offloading time of a layer’s activations. Therefore, full swapping with memory planning can achieve an MFU of 53.62% under 256K sequence length, far exceeding the 42.05% achieved by full recomputation with memory planning. However, for short sequence lengths, such as 64K, the offloading time of one layer’s activations block the GPU computation, resulting in a lower MFU than full recomputation. Full swapping presents another challenge as the sequence length grows longer: the host memory is rapidly depleted by offloaded activations, leading to OOHM errors.
By employing token-wise recomputation together with swapping, Memo consistently improves training efficiency for both short and long context lengths. For short sequence lengths like 64K, our tensor-level design only offloads the input tensor of the transformer layer and the FlashAttention output tensor to CPU memory, enabling efficient overlap of GPU computation and data transmission. For long context lengths, our token-level management successfully avoids depleting the CPU memory, and incurs only minimal recomputation overhead. Among all the methods, Memo supports the longest sequence length. Considering MFU, Memo achieves an average of $1.22\times$ MFU compared to full recomptutation with memory planning, and an average of $1.13\times$ MFU compared to full offloading with memory planning.
@@ 5.4. Scalability
To demonstrate the scalability of Memo, we train the 7B model on 8, 16, 32, and 64 GPUs respectively, and report the maximum supported sequence length. As shown in Figure 12(a), when the number of GPUs increases, the maximum sequence length supported by Memo increases linearly. When training on 8, 16, 32, and 64 GPUs, Memo is capable of training 1, 2, 4, 8 million sequence lengths, respectively, which demonstrates ideal scalability. Memo also consistently maintains an MFU of over 50% across different numbers of GPUs, as shown in Figure 12(b).
For DeepSpeed, as the number of GPUs increases, it can enlarge the SP size, leading to longer supported sequence lengths. Note that since because the 7B model has 32 attention heads and thereby the maximum SP size is 32, DeepSpeed achieves the same maximum sequence length 1536K for both 32 and 64 GPUs. Megatron-LM supports context parallelism, which has better scalability than DeepSpeed. When the number of GPUs increases, the longest sequence length it can handle grows sublinearly. Compared to the baselines, Memo introduces fine-grained activation memory management, achieving not only ideal scalability but also better efficiency.
We also evaluate the MFU metrics of the three systems when training the 7B model on 64 GPUs with sequence lengths varing from 1024K to 8192K. In Figure 12(c), as sequence length increases, the MFU of Memo maintains above 50% consistently, surpassing the competitors significantly.
5.5. Convergence of Memo
To demonstrate the correctness of our system implementation, we conduct a convergence experiment. Specifically, we train the 7B model with 128K sequence length on 8 GPUs and compare the convergence of Memo and Megatron-LM for 1000 iterations. For both systems, we fix the parallelism strategy to TP size 4 and CP size 2, and for Memo, we enumerate the value of $\alpha$ in $\{0,0.125,0.25,0.5,1\}$. As shown in Figure 12(d), the loss curves of Memo with different $\alpha$ values all align with Megatron-LM, confirming the correctness of our system implementation.
@@ 6. Related Work
Parallelism strategies for long context training: To tackle the challenge of long context training, DeepSpeed-Ulysses (Jacobs et al., 2023) employs an novel AllToAll communication to facilitate the partition of input sequence among GPUs, achieving lower communication overhead compared with Megatron-LM sequence parallelism (Korthikanti et al., 2022). LightSeq (Li et al., 2023), Ring Attention (Liu et al., 2023b), and Megatron-LM context parallelism (NVIDIA, 2024a) propose to split the sequence within self-attention computation, achieving better scalability. Recent efforts in the realm of sequence and context parallelisms (Chen et al., 2024; Fang and Zhao, 2024) aim to integrate multiple strategies and enhance existing distributed settings. It is worth noting that the fine-grained memory management of Memo is orthogonal to these distributed parallelism strategies.
Activation recomputation and swapping: Capuchin (Peng et al., 2020) proposes to combine recomputation and swapping to reduce memory footprint during training. The swapping decision is made by considering the tensor access pattern. In addition to tensor recomputation, MegTaichi (Hu et al., 2022) also proposes to co-optimize the tensor partition. Coop (Zhang et al., 2023) notices that naive tensor recomputation leads to severe memory fragmentation, and proposes heuristics to reduce memory fragmentation during tensor recomputation. While these works offer solutions for common deep learning models, they do not take advantage of the specific characteristics of LLM training to achieve full overlapping and fragmentation minimization.
Memory planning for deep learning models: The memory allocation problem of in deep learning models can be regarded as a DSA problem and solved by MIP (Sekiyama et al., 2018).OLLA (Steiner et al., 2022) proposes to optimize the lifetime and memory location of tensors during the training process by solving a joint ILP problem, reducing the peak memory during training. However, these methods does not exploit the repetitive substructure in LLMs and relies on heuristics to simplify the integer programming problem.
@@ 7. Conclusion
In this paper, we proposed Memo to address the memory challenges in long context LLM training. We designed a fine-grained activation recomputation and swapping strategy to fully utilize the idle PCIe bandwidth during the GPU computation, thereby reducing the activation rematerialization cost in long context LLM training. We employed a bi-level MIP technique to solve the problem of memory allocation within one transformer layer, and reused the same memory space for each identical layer so as to eliminate memory fragmentation. Through extensive experiments, we demonstrated that Memo achieved an average of $2.42\times$ MFU compared to Megatron-LM. By leveraging fine-grained tensor memory management, Memo achieved 52.30% MFU when training 7B LLM with 1 million sequence length on only 8 A800 GPUs.
Appendix A Detailed Parallelism Strategy in Evaluation
Table 56, and 7 present the detailed parallelism training strategies for end-to-end evaluation.
Table 5. Detailed parallelism strategies for DeepSpeed. “SP” stands for DeepSpeed-Ulysses sequence parallel size, “DP” stands for data parallel size, “ZeRO” stands for ZeRO stage, “AR” stands for activation recomputation.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
SP 2 4 8 oom oom oom oom oom oom oom oom oom
DP 4 2 1 oom oom oom oom oom oom oom oom oom
ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
8 7B AR On On On oom oom oom oom oom oom oom oom oom
SP 4 8 8 oom oom oom oom oom oom oom oom oom
DP 4 2 2 oom oom oom oom oom oom oom oom oom
ZeRO 3 3 3 oom oom oom oom oom oom oom oom oom
16 13B AR On On On oom oom oom oom oom oom oom oom oom
SP 8 8 oom oom oom oom oom oom oom oom oom oom
DP 4 4 oom oom oom oom oom oom oom oom oom oom
ZeRO 3 3 oom oom oom oom oom oom oom oom oom oom
32 30B AR On On oom oom oom oom oom oom oom oom oom oom
SP 8 8 16 32 32 32 64 64 64 64 64 oom
DP 8 8 4 2 2 2 1 1 1 1 1 oom
ZeRO 3 3 3 3 3 3 3 3 3 3 3 oom
64 65B AR On On On On On On On On On On On oom
Table 6. Detailed parallelism strategies for Megatron-LM. “TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, “AR” stands for activation recomputation. For all evaluations of Megatron-LM, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
TP 2 4 4 4 oom oom oom oom oom oom oom oom
CP 4 2 2 2 oom oom oom oom oom oom oom oom
DP 1 1 1 1 oom oom oom oom oom oom oom oom
$P$ 1 1 1 1 oom oom oom oom oom oom oom oom
8 7B AR On On On On oom oom oom oom oom oom oom oom
TP 8 4 4 4 8 8 oom oom oom oom oom oom
CP 2 2 2 2 2 2 oom oom oom oom oom oom
DP 1 1 1 1 1 1 oom oom oom oom oom oom
$P$ 1 2 2 2 1 1 oom oom oom oom oom oom
16 13B AR On On On On On On oom oom oom oom oom oom
TP 4 4 8 8 oom oom oom oom oom oom oom oom
CP 4 8 4 4 oom oom oom oom oom oom oom oom
DP 1 1 1 1 oom oom oom oom oom oom oom oom
$P$ 2 1 1 1 oom oom oom oom oom oom oom oom
32 30B AR On On On On oom oom oom oom oom oom oom oom
TP 8 32 16 16 64 oom oom oom oom oom oom oom
CP 8 1 1 4 1 oom oom oom oom oom oom oom
DP 1 2 1 1 1 oom oom oom oom oom oom oom
$P$ 1 1 4 1 1 oom oom oom oom oom oom oom
64 65B AR On On On On On oom oom oom oom oom oom oom
Table 7. Detailed parallelism strategies for Memo. “TP” stands for tensor parallel size, “CP” stands for context parallel size, “DP” stands for data parallel size, “PP” stands for pipeline parallel size, $\alpha$ stands for CPU offload fraction. For all evaluations of Memo, Megatron-style sequence parallel and ZeRO-1 optimizer are enabled.
Sequence Length
# GPUs Model Size Strategy 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
TP 2 4 4 4 4 4 4 4 8 oom oom oom
CP 2 2 2 2 2 2 2 2 1 oom oom oom
DP 2 1 1 1 1 1 1 1 1 oom oom oom
$P$ 1 1 1 1 1 1 1 1 1 oom oom oom
8 7B $\alpha$ 0.0 0.0 1.0 0.5 0.125 0.125 0.0 0.0 0.0 oom oom oom
TP 2 4 4 4 4 4 4 4 4 4 4 4
CP 2 2 2 2 4 4 4 4 4 4 4 4
DP 4 2 2 1 1 1 1 1 1 1 1 1
$P$ 1 1 1 2 1 1 1 1 1 1 1 1
16 13B $\alpha$ 0.0 0.0 0.0 0.0 0.5 0.125 0.125 0.0 0.0 0.0 0.0 0.0
TP 4 8 8 8 8 8 8 8 8 8 8 oohm
CP 1 1 2 2 2 4 4 4 4 4 4 oohm
DP 8 4 2 2 2 1 1 1 1 1 1 oohm
$P$ 1 1 1 1 1 1 1 1 1 1 1 oohm
32 30B $\alpha$ 0.0 0.0 0.0 0.0 0.0 0.25 0.125 0.125 0.125 0.125 0.0 oohm
TP 8 4 8 8 8 8 8 8 8 8 8 8
CP 1 4 4 8 8 8 8 8 8 8 8 8
DP 8 4 2 1 1 1 1 1 1 1 1 1
$P$ 1 1 1 1 1 1 1 1 1 1 1 1
64 65B AR 0.0 0.0 0.0 0.25 0.25 0.25 0.25 0.125 0.125 0.0 0.0 0.0

### v2

@@ 5.4. Scalability
To show the scalability of Memo, we train the 7B model on 8, 16, 32, and 64 GPUs respectively, and report the maximum supported sequence length. As shown in Figure 11(a), when the number of GPUs increases, the maximum sequence length supported by Memo increases linearly. When training on 8, 16, 32, and 64 GPUs, Memo is capable of training 1, 2, 4, 8 million sequence lengths, respectively, which demonstrates ideal scalability. Memo also consistently maintains an MFU of over 50% across different numbers of GPUs, as shown in Figure 11(b).
For DeepSpeed, as the number of GPUs increases, it can enlarge the SP degree, leading to longer supported sequence length. Note that since the 7B model has 32 attention heads, the maximum SP degree is 32. As a result, DeepSpeed achieves the same maximum sequence length of 1536K for both 32 and 64 GPUs. Megatron-LM supports context parallelism, which has better scalability. When the number of GPUs increases, the longest sequence length it can handle grows sublinearly. Compared to the baselines, Memo introduces fine-grained activation memory management, achieving not only ideal scalability but also better efficiency.
We also evaluate the MFU of the three systems when training the 7B model on 64 GPUs with sequence lengths varing from 1024K to 8192K. In Figure 11(c), as sequence length increases, the MFU of Memo maintains above 50%, surpassing the baselines.
$5.$ Convergence of Memo
To confirm the correctness of our system implementation, we conduct a convergence experiment. We train a 7B model with 128K sequence length on 8 GPUs and compare the convergence of Memo and Megatron-LM for 1000 iterations. For both systems, we fix the parallelism strategy to TP=4=4 and CP=2=2. For Memo, we enumerate the value of $\alpha$ in $\{0,0.125,0.25,0.5,1\}$. As shown in Figure 11(d), the loss curves of Memo with different $\alpha$ values all align with Megatron-LM, confirming the correctness of Memo.
@@ 6. Related Work
Data management and machine learning: Recently, the intersection of data management and machine learning has gained popularity, with numerous studies emerging from the data management community (Miao et al., 2024a, 2023; Zhang et al., 2022; Zhao et al., 2023; Murray et al., 2021; Um et al., 2023; Isenko et al., 2022; Mohan et al., 2021; Miao et al., 2021; Li et al., 2023b; Gao et al., 2024; Miao et al., 2022; Wang et al., 2024; Yuan et al., 2021; Khamis et al., 2020; Nakandala et al., 2020; Koutsoukos et al., 2021; Schleich et al., 2023; Zhao et al., 2022; Zhou et al., 2024; Wang et al., 2023; Yu et al., 2023; Ooi et al., 2024; Zhang et al., 2023b, 2024a). Some works focus on optimizing I/O data management to accelerate data pre-processing pipelines (Murray et al., 2021; Um et al., 2023; Isenko et al., 2022; Mohan et al., 2021). Another work proposes to use a fine-grained automatic parallelism strategy to speed up model training (Miao et al., 2022). Memo also employs data management techniques and methodologies to enhance the efficiency of machine learning tasks, specifically addressing the challenges associated with long-sequence LLM training.
Parallelism strategies for long context training: To tackle the challenge of long context training, DeepSpeed-Ulysses (Jacobs et al., 2023) employs an novel AllToAll communication to facilitate the partition of input sequence among GPUs, achieving lower communication overhead compared with Megatron-LM sequence parallelism (Korthikanti et al., 2022). LightSeq (Li et al., 2023a), Ring Attention (Liu et al., 2023), and Megatron-LM context parallelism (NVIDIA, 2024a) propose to split the sequence within self-attention computation, achieving better scalability. Recent efforts in the realm of sequence and context parallelisms (Chen et al., 2024; Fang and Zhao, 2024; Ge et al., 2024) aim to integrate multiple strategies and enhance existing distributed settings. It is worth noting that the fine-grained memory management of Memo is orthogonal to these distributed parallelism strategies.
Activation recomputation and swapping: Capuchin (Peng et al., 2020) proposes to combine recomputation and swapping to reduce memory footprint during training. The swapping decision is made by considering the tensor access pattern. In addition to tensor recomputation, MegTaichi (Hu et al., 2022) also proposes to co-optimize the tensor partition. Coop (Zhang et al., 2023a) notices that naive tensor recomputation leads to severe memory fragmentation, and proposes heuristics to reduce memory fragmentation during tensor recomputation. While these works offer solutions for common deep learning models, they do not take advantage of the specific characteristics of LLM training to achieve full overlapping and fragmentation minimization.
Memory planning for deep learning models: The memory allocation problem in deep learning models can be regarded as a DSA problem and solved by MIP (Sekiyama et al., 2018). OLLA (Steiner et al., 2022) proposes to optimize the lifetime and memory location of tensors during the training process by solving a joint ILP problem, reducing the peak memory during training. However, it does not exploit the repetitive substructure in LLMs and relies on heuristics to simplify the integer programming problem.
@@ 7. Conclusion
In this paper, we proposed Memo to address the memory challenges in long context LLM training. We designed a fine-grained activation recomputation and swapping strategy to fully utilize the idle PCIe bandwidth during the GPU computation, thereby reducing the activation rematerialization cost in long context LLM training. We employed a bi-level MIP technique to first solve the problem of memory allocation within one transformer layer, and then reused the same memory space for each identical layer so as to eliminate memory fragmentation. Through extensive experiments, we demonstrated that Memo achieved an average of $1.97\times$ MFU compared to Megatron-LM. By leveraging fine-grained tensor memory management, Memo achieved 52.30% MFU when training 7B LLM with 1 million sequence length on only 8 A800 GPUs.
Acknowledgement
This work is supported by National Science and Technology Major Project (2022ZD0116315), National Natural Science Foundation of China (U22B2037, U23B2048, 62402011), Beijing Municipal Science and Technology Project (Z231100010323002), China National Postdoctoral Program for Innovative Talents (BX20230012), China Postdoctoral Science Foundation (2024M750103), Beijing Natural Science Foundation (4244080), research grant No. IPT-2024JK29, PKU-Tencent joint research Lab, and High-performance Computing Platform of Peking University. Fangcheng Fu and Bin Cui are the corresponding authors.